Find the demensions of a rectangle a with the greatest area whose perimeter is 30 feet.
Answer by checkley77(12569)

30/4=7.5 for each side of a square.
7.5*7.5=56.25 maximum area for the square.
Proof for all other dimentions:
7.4*7.6=56.24
7*8=56
10*5=50